# Attenuation Bias

### v2

@@ A.IV.1.5 American Politics: Ideology & Presidential Vote
We here consider an application from American politics, in particular, from the literature on the relationship between latent ideology and partisan vote choice—a relationship that has been extensively studied and shown to be quite strong [26, 46]. To what extent might measurement error in the estimation of latent ideology lead to over- or under-estimation in estimated relationships with the various strategies previously described?
Using the same design as outlined in Section 7.1, we explore the relationship between latent ideology, as measured by 30 questions on the 2016 CCES regarding respondents’ policy positions, and vote choice for president. We again vary the number of observations used and policy factors considered in estimating latent traits, using random sub-samples and computing average estimates. Latent factors are estimated as described above using the expectation-maximization algorithm.
@@ A.IV.1.8 Additional Information, Knowledge Example
Appendix V: A lpme Package Tutorial
lpme implements a bootstrap analysis for latent variable models with measurement error correction. We average over nPartition split sample partitions:
@@
devtools::install_github("cjerzak/lpme/lpme-software")
# Generate data
Yobs <- rnorm(1000)
ObservablesMat <- matrix(sample(c(0, 1), 1000 * 10,
                         replace = TRUE),
                         ncol = 10)
@@
results <- lpme::lpme(
  Yobs,
  ObservablesMat,
  nBoot = 32L,
  nPartition = 10L,
  bootBasis = 1:length(Yobs),
  ReturnIntermediaries = TRUE
# Print the corrected IV coefficient and its standard error
print(c(results\$Corrected_IVRegCoef, results\$Corrected_IVRegSE))

### v3

@@ A.IV.1.5 American Politics: Ideology & Presidential Vote
We here consider an application from American politics, in particular, from the literature on the relationship between latent ideology and partisan vote choice—a relationship that has been extensively studied and shown to be quite strong [undefv, undefan]. To what extent might measurement error in the estimation of latent ideology lead to over- or under-estimation in estimated relationships with the various strategies previously described?
We explore the relationship between latent ideology, as measured by 30 questions on the 2016 CCES regarding respondents’ policy positions, and vote choice for president. We again vary the number of observations used and policy factors considered in estimating latent traits, using random sub-samples and computing average estimates. Latent factors are estimated as described above using the expectation-maximization algorithm.
@@ A.IV.1.7 Effects of Political Knowledge
Figure A.IV.3: Left. Scatterplot of standardized sentiment scores computed from the first (x-axis) and second (y-axis) halves of each text. Points represent individual observations; the dashed diagonal line indicates perfect agreement; the solid line shows the linear regression fit with prediction intervals. Right. Coefficient by decade.
@@ A.IV.1.8 Additional Information, Knowledge Example
Appendix V: A lpmec Package Tutorial
lpmec is an R package for latent-predictor measurement error correction under identification restrictions (i.e., when a key regressor is an estimated latent trait from observed indicators, leading to attenuation bias in outcome regressions). Its core workflow repeatedly splits the indicator matrix into independent halves (and can average over many such partitions), estimates the latent score separately in each half (via EM IRT for binary/ordinal items using emIRT, PCA, simple averaging, or a user-supplied estimator), and then applies split-sample IV and OLS-style corrections that use the cross-half relationship to de-attenuate regression coefficients. The main entry points are lpmec_onerun() (single partition) and lpmec() (bootstrap + partition aggregation with percentile confidence intervals), with S3 print/summary/plot methods and built-in data (e.g., KnowledgeVoteDuty); the codebase also includes sensitivity reporting via sensemakr’s extreme robustness value.
For Bayesian estimation, lpmec supports MCMC either through an R backend (pscl::ideal) or an optional Python backend (numpyro via reticulate), plus an over-imputation variant using Amelia. The mcmc_joint option implements a joint Bayesian model that couples an IRT measurement model for the indicators with an outcome model (e.g., $Y_{i}\sim\mathcal{N}(\alpha+\beta\theta_{i},\sigma)$), so posterior uncertainty in the latent trait propagates directly into inference on the regression slope. When using the NumPyro/JAX backend, this can take advantage of JAX’s JIT/XLA compilation (and, optionally, GPU/TPU acceleration) and supports subsampling-based kernels for larger problems.
We average over n_partition split sample partitions:
@@
devtools::install_github(
  "cjerzak/lpmec-software",
  subdir = "lpmec"
library(lpmec)
data(KnowledgeVoteDuty)
@@
results <- lpmec::lpmec(
  Y = KnowledgeVoteDuty\$voteduty,
  observables = as.matrix(KnowledgeVoteDuty[, 2:5]),
  n_boot = 32L,
  n_partition = 10L,
  estimation_method = "em"
print(results)
summary(results)